# definition

### v2

@@ D.3 Properties of $h_{\textnormal{ID}}(\cdot,\cdot)$
$\leq\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left(h({\bm{X}}_{t+1},[Nm^{\prime}])-h({\bm{X}}_{t}P,[Nm^{\prime}])\right)$
$\leq\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left\lvert\sum_{s\in\mathbb{S}}\sum_{i\in[Nm^{\prime}]}\epsilon_{i,t}(s)\sum_{s^{\prime}\in\mathbb{S}}\left(g_{i}(s^{\prime})\left(\frac{P_{i}-\Xi_{i}}{\gamma}\right)^{\ell}(s,s^{\prime})\right)\right\rvert\,,$ (52)
where $\epsilon_{i,t}$ is a $\lvert\mathbb{S}\rvert$-dimensional random vector given by $\epsilon_{i,t}\triangleq X_{i,t+1}-X_{i,t}P_{i}$, which satisfies $\mathbb{E}\left[\epsilon_{i,t}(s)\;\middle|\;{\bm{X}}_{t}\right]=0$ and $\left\lvert\epsilon_{i,t}(s)\right\rvert\leq 1$ for any $s\in\mathbb{S}$. Now, we take the expectations of the positive parts of the inequality (52) conditioned on $X_{t}$. We deduce that
@@
$\leq\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\left\lvert\sum_{s\in\mathbb{S}}\sum_{i\in[Nm^{\prime}]}\epsilon_{i,t}(s)\sum_{s^{\prime}\in\mathbb{S}}\left(g_{i}(s^{\prime})\left(\frac{P_{i}-\Xi_{i}}{\gamma}\right)^{\ell}(s,s^{\prime})\right)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\sum_{s\in\mathbb{S}}\left\lvert\sum_{i\in[Nm^{\prime}]}\epsilon_{i,t}(s)\sum_{s^{\prime}\in\mathbb{S}}\left(g_{i}(s^{\prime})\left(\frac{P_{i}-\Xi_{i}}{\gamma}\right)^{\ell}(s,s^{\prime})\right)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\sum_{s\in\mathbb{S}}\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left\lvert\sum_{i\in[Nm^{\prime}]}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]\,,$ (53)
where $q_{g,i}^{\ell}(s)\triangleq\sum_{s^{\prime}}g_{i}(s^{\prime})\left(\frac{P_{i}-\Xi_{i}}{\gamma}\right)^{\ell}(s,s^{\prime})$ for any $\ell\in\mathbb{N}_{+}$, $g\in\mathcal{G}$, $i\in[Nm]$ and $s\in\mathbb{S}$. To bound the summand, obverve that for each $t$ and $s$, $\{\epsilon_{i,t}(s)\}_{i\in[Nm]}$ are independent and have zero means. Consequently,
$\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left\lvert\sum_{i\in[Nm^{\prime}]}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\mathbb{E}\left[\left(\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left\lvert\sum_{i\in[Nm^{\prime}]}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right\rvert\right)^{2}\;\middle|\;{\bm{X}}_{t}\right]^{1/2}$ (54)
$\leq 2\mathbb{E}\left[\sum_{i\in[Nm]}\left(q_{g,i}^{\ell}(s)\right)^{2}\right]^{1/2}$ (55)
$\leq 2\sqrt{N}\max_{i\in[Nm]}\left\lvert q_{g,i}^{\ell}(s)\right\rvert\,,$
where (54) is due to Cauchy–Schwarz; in (55), we apply Doob’s $L_{2}$ inequality to the submartingale $\left\{\left\lvert\sum_{i\in[n]}\epsilon_{i,t}(s)q_{g,i}^{\ell}(s)\right\rvert\right\}_{n=1}^{Nm}$. Finally, following similar arguments as those in the proof of Lemma 3, we get
$\leq 2\sqrt{N}\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\sum_{s\in\mathbb{S}}\max_{i\in[Nm]}\left\lvert q_{g,i}^{\ell}(s)\right\rvert$
$\leq 2C_{h}\sqrt{N}.$
In particular, the last inequality follows from the same calculations as those in (46) and (47). ∎
Appendix E Lemmas for focus set
@@
where $\widehat{A}_{i,t}$’s are the actions sampled from the optimal single-armed policies by the ID policy.
Below we state the three lemmas, and we then prove them in the subsections.
@@
for some constant $K_{\text{conf}}>0$.
@@
for some constant $K_{\text{mono}}>0$.
@@
for some constant $K_{\text{cov}}>0$.
@@
E.1 Proof of Lemma 10 (Majority conformity)
@@
$\Delta_{t}=\max_{k\in[K]}\max_{m\in[0,1]_{N}}\left\lvert\sum_{i\in[Nm]}\left(c_{k,i}(S_{i,t},\widehat{A}_{i,t})-\sum_{s}X_{i,t}(s)c_{k,i}^{*}(s)\right)\right\rvert\,,$ (59)
@@
$\leq h_{\textnormal{ID}}({\bm{X}}_{t},n/N)+\max_{k\in[K]}\left\lvert\sum_{i\in[n]}\left(c_{k,i}(S_{i,t},\widehat{A}_{i,t})-\sum_{s}X_{i,t}(s)c_{k,i}^{*}(s)\right)\right\rvert$ (61)
$\leq h_{\textnormal{ID}}({\bm{X}}_{t},n/N)+\Delta_{t}.$ (62)
where (61) inequality is because
@@
$=\min_{k}\overline{C}^{*}_{k}([Nm])-\min_{k}\overline{C}^{*}_{k}([Nm({\bm{X}}_{t})])$
$\geq\min_{k}\left(\overline{C}^{*}_{k}([Nm])-\overline{C}^{*}_{k}([Nm({\bm{X}}_{t})])\right)$
@@
where (E.1) follows from the strict slope of the slack budget $\overline{C}^{*}([Nm])$ (Lemma 2). Therefore, choosing $m=m({\bm{X}}_{t})-(\Delta_{t}+M_{c})/(\eta_{c}N)$, we obtain
@@
It remains to upper bound the conditional expectation of $\Delta_{t}$ given ${\bm{X}}_{t}$. Define the random variable $\xi_{k,i}\triangleq c_{k,i}(S_{i,t},\widehat{A}_{i,t})-\sum_{s}X_{i,t}(s)c_{k,i}^{*}(s)$ for each arm $i\in[N]$ and cost type $k\in[K]$. Subsequently, $\mathbb{E}\left[\Delta_{t}\;\middle|\;{\bm{X}}_{t}\right]$ can be rewritten and bounded as
@@
where (65) follows from the Cauchy–Schwarz inequality. Next, we argue that the sequence $\left\{\left\lvert\sum_{i\in[n]}\xi_{k,i}\right\rvert\right\}_{n\in[N]}$ is a submartingale, which enables us to apply Doob’s $L_{2}$ inequality (Durrett, 2019, Theorem 5.4.3) to bound the expression in (65). Observe that, for each cost type $k$, conditioned on ${\bm{X}}_{t}$, the sequence of random variables $\left\{\xi_{k,i}\right\}_{i\in[N]}$ are independent and have zero conditional means. Consequently, the sequence of partial sums $\left\{\sum_{i\in[n]}\xi_{k,i}\right\}_{n\in[N]}$ forms a martingale, which becomes a submartingale upon applying taking absolute values. Thus, by applying Doob’s $L_{2}$ inequality and utilizing the bound $\xi_{k,i}\leq c_{\max}$, we obtain
$\mathbb{E}\left[\Delta_{t}\;\middle|\;{\bm{X}}_{t}\right]\leq\sum_{k\in[K]}2\mathbb{E}\left[\sum_{i\in[N]}\xi_{k,i}^{2}\;\middle|\;{\bm{X}}_{t}\right]^{1/2}\leq 2Kc_{\max}\sqrt{N}\,.$ (66)
Combining (66) with the previous calculations, and taking $K_{\text{conf}}\triangleq(2Kc_{\max}+M_{c})/\eta_{c}$, we conclude that
@@
E.2 Proof of Lemma 11 (Almost non-shrinking)
@@
$\mathbb{E}\left[\big{(}h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t})\big{)}^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq 2\big{(}C_{h}+\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}\big{)}\sqrt{N}\,,$
@@
$\leq\mathbb{E}\left[\sum_{i\in[Nm({\bm{X}}_{t})]\backslash[N_{t}^{*}]}\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\left\lVert X_{i,t+1}-X_{i,t+1}^{\prime}\right\rVert\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert\gamma^{-\ell}\left\lVert g_{i}\right\rVert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\mathbb{E}\left[\sum_{i\in[Nm({\bm{X}}_{t})]\backslash[N_{t}^{*}]}2\lvert\mathbb{S}\rvert^{1/2}\sum_{\ell=0}^{\infty}\gamma^{-\ell}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert\sum_{g\in\mathcal{G}}\left\lVert g_{i}\right\rVert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\mathbb{E}\left[\sum_{i\in[Nm({\bm{X}}_{t})]\backslash[N_{t}^{*}]}2\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})\;\middle|\;{\bm{X}}_{t}\right]$
$\leq 2\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})\mathbb{E}\left[(Nm({\bm{X}}_{t})-N_{t}^{*})^{+}\;\middle|\;{\bm{X}}_{t}\right]$ (68)
$\leq 2\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}\sqrt{N}\,,$ (69)
@@
$\mathbb{E}\left[\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq 2\left(C_{h}+\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}\right)\sqrt{N}\,.$
@@
$m({\bm{X}}_{t+1})\geq m({\bm{X}}_{t})-\frac{1}{\eta_{c}N}\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}-\frac{M_{c}}{\eta_{c}N}$ (70)
To prove the claim, by the maximality of $m({\bm{X}}_{t+1})$, it suffices to show that for any $\bar{m}\in[0,1]_{N}$ such that
$\bar{m}\leq m({\bm{X}}_{t})-\frac{1}{\eta_{c}N}\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}-\frac{M_{c}}{\eta_{c}N}$ (71)
@@
$\mathbb{E}\left[(m({\bm{X}}_{t})-m({\bm{X}}_{t+1}))^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq\frac{1}{N}\mathbb{E}\left[\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]+\frac{M_{c}}{N\eta_{c}},$
where the right-hand side can be further bounded using Lemma 13 which states that
$\mathbb{E}\left[\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq 2\left(C_{h}+\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}\right)\sqrt{N}\,.$
Therefore, we have:
$\mathbb{E}\left[(m({\bm{X}}_{t})-m({\bm{X}}_{t+1}))^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq 2\left(C_{h}+\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}\right)\frac{1}{\sqrt{N}}+\frac{M_{c}}{\eta_{c}N},$
@@
with $K_{\text{mono}}\triangleq 2\left(C_{h}+\lvert\mathbb{S}\rvert^{1/2}C_{\gamma}(Kc_{\max}+r_{\max})K_{\text{conf}}\right)+M_{c}/\eta_{c}$. ∎
E.3 Proof of Lemma 12 (Sufficient coverage)

### v3

@@ E.3 Properties of $h_{\textnormal{ID}}(\cdot,\cdot)$
$\leq\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left(h({\bm{X}}_{t+1},[Nm^{\prime}])-h({\bm{X}}_{t}P,[Nm^{\prime}])\right)$ (49)
$\leq\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\gamma^{-\ell}\left\lvert\sum_{i\in[Nm^{\prime}]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert\,,$ (50)
where (49) follows from (39); $\epsilon_{i,t}$ is a $\lvert\mathbb{S}\rvert$-dimensional random vector given by $\epsilon_{i,t}\triangleq X_{i,t+1}-X_{i,t}P_{i}$, which satisfies $\mathbb{E}\left[\epsilon_{i,t}(s)\;\middle|\;{\bm{X}}_{t}\right]=0$ for any $s\in\mathbb{S}$ and $\left\lVert\epsilon_{i,t}\right\rVert_{1}\leq 2$; (50) follows (40). Now, we take the expectations of the positive parts of the inequality (50) conditioned on $X_{t}$, which yields
@@
$\leq\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\max_{g\in\mathcal{G}}\sup_{\ell\in\mathbb{N}}\gamma^{-\ell}\left\lvert\sum_{i\in[Nm^{\prime}]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\sum_{\ell=0}^{\infty}\gamma^{-\ell}\sum_{g\in\mathcal{G}}\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left\lvert\sum_{i\in[Nm^{\prime}]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\sum_{\ell=0}^{\infty}\gamma^{-\ell}\sum_{g\in\mathcal{G}}\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left\lvert\sum_{i\in[Nm^{\prime}]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert^{2}\;\middle|\;{\bm{X}}_{t}\right]^{1/2}.$ (51)
Because $\epsilon_{i,t}$ are independent across $i\in[Nm]$ and $\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle$ has zero mean conditioned on ${\bm{X}}_{t}$, the partial sum $\sum_{i\in[n]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle$ is a martingale in $n$ and $\left\lvert\sum_{i\in[n]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert$ is a submartingale in $n$ for $n\in[Nm]$. Then it follows from Doob’s $L_{2}$ inequality (11, Theorem 5.4.3) that
$\mathbb{E}\left[\max_{m^{\prime}\in[0,1]_{N}\colon m^{\prime}\leq m}\left\lvert\sum_{i\in[Nm^{\prime}]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\right\rvert^{2}\;\middle|\;{\bm{X}}_{t}\right]^{1/2}$
$\leq 2\mathbb{E}\left[\bigg{(}\sum_{i\in[Nm]}\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle\bigg{)}^{2}\;\middle|\;{\bm{X}}_{t}\right]^{1/2}$
$=2\left(\sum_{i\in[Nm]}\mathbb{E}\left[\left\langle\epsilon_{i,t},(P_{i}-\Xi_{i})^{\ell}g_{i}\right\rangle^{2}\;\middle|\;{\bm{X}}_{t}\right]\right)^{1/2}$
$\leq 2\left(\sum_{i\in[Nm]}\mathbb{E}\left[\Big{(}\left\lVert\epsilon_{i,t}\right\rVert_{1}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}\left\lVert g_{i}\right\rVert_{\infty}\Big{)}^{2}\;\middle|\;{\bm{X}}_{t}\right]\right)^{1/2}$
$\leq 4\sqrt{Nm}\max_{i\in[Nm]}\left\lVert g_{i}\right\rVert_{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty},$ (52)
where (52) uses the fact that $\left\lVert\epsilon_{i,t}\right\rVert_{1}\leq 2$. Substituting the bound in (52) back to (51), we get
$\mathbb{E}\left[(h_{\textnormal{ID}}({\bm{X}}_{t+1},m)-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m))^{+}\;\middle|\;{\bm{X}}_{t}\right]$ $\leq 4\sqrt{Nm}\sum_{\ell=0}^{\infty}\gamma^{-\ell}\sum_{g\in\mathcal{G}}\max_{i\in[Nm]}\left\lVert g_{i}\right\rVert_{\infty}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}$
$=4\sqrt{N}(Kc_{\max}+r_{\max})\sum_{\ell=0}^{\infty}\gamma^{-\ell}\max_{i\in[N]}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}$
$\leq 4\sqrt{N}(Kc_{\max}+r_{\max})C_{\tau}$
$=2C_{h}\sqrt{N}.\qed$
Appendix F Lemmas for focus set
@@
where $\widehat{A}_{i,t}$’s are the ideal actions sampled from the optimal single-armed policies by the ID policy.
Below we state the three lemmas, and we prove them in the subsections.
@@
for some constant $K_{\text{conf}}>0$ independent of $N$.
@@
for some constant $K_{\text{mono}}>0$ independent of $N$.
@@
for some constant $K_{\text{cov}}>0$ independent of $N$.
@@
F.1 Proof of Lemma 8 (Majority conformity)
@@
$\Delta_{t}=\max_{k\in[K]}\max_{m\in[0,1]_{N}}\left\lvert\sum_{i\in[Nm]}\left(c_{k,i}(S_{i,t},\widehat{A}_{i,t})-\left\langle X_{i,t},c_{k,i}^{*}\right\rangle\right)\right\rvert\,,$ (56)
@@
$\leq h_{\textnormal{ID}}({\bm{X}}_{t},n/N)+\max_{k\in[K]}\left\lvert\sum_{i\in[n]}\left(c_{k,i}(S_{i,t},\widehat{A}_{i,t})-\left\langle X_{i,t},c_{k,i}^{*}\right\rangle\right)\right\rvert$ (58)
$\leq h_{\textnormal{ID}}({\bm{X}}_{t},n/N)+\Delta_{t},$ (59)
where the inequality (58) follows from the argument that
@@
$=\min_{k\in[K]}\overline{C}^{*}_{k}([Nm])-\min_{k\in[K]}\overline{C}^{*}_{k}([Nm({\bm{X}}_{t})])$
$\geq\min_{k\in[K]}\left(\overline{C}^{*}_{k}([Nm])-\overline{C}^{*}_{k}([Nm({\bm{X}}_{t})])\right)$
@@
where (F.1) follows from the strict slope of the remaining budget $\overline{C}^{*}([Nm])$ (Lemma 2). Therefore, choosing $m=m({\bm{X}}_{t})-(\Delta_{t}+M_{c})/(\eta_{c}N)$, we obtain
@@
It remains to upper bound the conditional expectation of $\Delta_{t}$ given ${\bm{X}}_{t}$. We define the random variable $\xi_{k,i}\triangleq c_{k,i}(S_{i,t},\widehat{A}_{i,t})-\langle X_{i,t},c_{k,i}^{*}\rangle$ for each arm $i\in[N]$ and cost type $k\in[K]$. Subsequently, $\mathbb{E}\left[\Delta_{t}\;\middle|\;{\bm{X}}_{t}\right]$ can be rewritten and bounded as
@@
where (62) follows from the Cauchy–Schwarz inequality. Next, we argue that the sequence $\left\{\left\lvert\sum_{i\in[n]}\xi_{k,i}\right\rvert\right\}_{n\in[N]}$ is a submartingale, which allows us to apply Doob’s $L_{2}$ inequality (11, Theorem 5.4.3) to bound the expression in (62). Observe that, for each cost type $k$, conditioned on ${\bm{X}}_{t}$, the sequence of random variables $\left\{\xi_{k,i}\right\}_{i\in[N]}$ are independent and have zero conditional means. Consequently, the sequence of partial sums $\left\{\sum_{i\in[n]}\xi_{k,i}\right\}_{n\in[N]}$ forms a martingale, which becomes a submartingale upon taking the absolute value. Thus, by applying Doob’s $L_{2}$ inequality and utilizing the bound $\xi_{k,i}\leq c_{\max}$, we obtain
$\mathbb{E}\left[\Delta_{t}\;\middle|\;{\bm{X}}_{t}\right]$ $\leq\sum_{k\in[K]}2\mathbb{E}\left[\bigg{(}\sum_{i\in[N]}\xi_{k,i}\bigg{)}^{2}\;\middle|\;{\bm{X}}_{t}\right]^{1/2}$
$=\sum_{k\in[K]}2\bigg{(}\sum_{i\in[N]}\mathbb{E}\left[\xi_{k,i}^{2}\;\middle|\;{\bm{X}}_{t}\right]\bigg{)}^{1/2}$
$\leq 2Kc_{\max}\sqrt{N}\,.$ (63)
Combining (63) with our earlier calculations, we get
@@
where $K_{\text{conf}}\triangleq(2Kc_{\max}+M_{c})/\eta_{c}$. ∎
F.2 Proof of Lemma 9 (Almost non-shrinking)
@@
$\mathbb{E}\left[\big{(}h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t})\big{)}^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq\big{(}2+K_{\text{conf}}\big{)}C_{h}\sqrt{N}\,,$
@@
$=\mathbb{E}\left[\sum_{i\in[Nm({\bm{X}}_{t})]\backslash[N_{t}^{*}]}\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\left\lvert\sum_{i\in[Nm^{\prime}]}\left\langle(X_{i,t+1}^{\prime}-X_{i,t+1})(P_{i}-\Xi_{i})^{\ell}\gamma^{-\ell},g_{i}\right\rangle\right\rvert\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\mathbb{E}\left[\sum_{i\in[Nm({\bm{X}}_{t})]\backslash[N_{t}^{*}]}\sum_{g\in\mathcal{G}}\sum_{\ell=0}^{\infty}\gamma^{-\ell}\left\lVert X_{i,t+1}-X_{i,t+1}^{\prime}\right\rVert_{1}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}\left\lVert g_{i}\right\rVert_{\infty}\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\mathbb{E}\left[\sum_{i\in[Nm({\bm{X}}_{t})]\backslash[N_{t}^{*}]}2\sum_{\ell=0}^{\infty}\gamma^{-\ell}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}\sum_{g\in\mathcal{G}}\left\lVert g_{i}\right\rVert_{\infty}\;\middle|\;{\bm{X}}_{t}\right]$
$\leq\mathbb{E}\left[\sum_{i\in[Nm({\bm{X}}_{t})]\backslash[N_{t}^{*}]}2C_{\tau}(Kc_{\max}+r_{\max})\;\middle|\;{\bm{X}}_{t}\right]$
$\leq 2C_{\tau}(Kc_{\max}+r_{\max})\mathbb{E}\left[(Nm({\bm{X}}_{t})-N_{t}^{*})^{+}\;\middle|\;{\bm{X}}_{t}\right]$ (65)
$\leq 2C_{\tau}(Kc_{\max}+r_{\max})K_{\text{conf}}\sqrt{N}\,,$ (66)
@@
$\mathbb{E}\left[\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-\gamma h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]$ $\leq\left(2C_{h}+2C_{\tau}(Kc_{\max}+r_{\max})K_{\text{conf}}\right)\sqrt{N}$
$=\left(2C_{h}+C_{h}K_{\text{conf}}\right)\sqrt{N},$
where the last equality is because $C_{h}=2C_{\tau}(Kc_{\max}+r_{\max})$. ∎
@@
$m({\bm{X}}_{t+1})\geq m({\bm{X}}_{t})-\frac{1}{\eta_{c}N}\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}-\frac{M_{c}}{\eta_{c}N}.$ (67)
To prove the claim, by the maximality of $m({\bm{X}}_{t+1})$ it suffices to show that for any $\bar{m}\in[0,1]_{N}$ s.t.
$\bar{m}\leq m({\bm{X}}_{t})-\frac{1}{\eta_{c}N}\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}-\frac{M_{c}}{\eta_{c}N},$ (68)
@@
$\mathbb{E}\left[(m({\bm{X}}_{t})-m({\bm{X}}_{t+1}))^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq\frac{1}{\eta_{c}N}\mathbb{E}\left[\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]+\frac{M_{c}}{N\eta_{c}},$
where the right-hand side can be further bounded using Lemma 11, which states that
$\mathbb{E}\left[\left(h_{\textnormal{ID}}({\bm{X}}_{t+1},m({\bm{X}}_{t}))-h_{\textnormal{ID}}({\bm{X}}_{t},m({\bm{X}}_{t}))\right)^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq\big{(}2+K_{\text{conf}}\big{)}C_{h}\sqrt{N}\,.$
Therefore, we have
$\mathbb{E}\left[(m({\bm{X}}_{t})-m({\bm{X}}_{t+1}))^{+}\;\middle|\;{\bm{X}}_{t}\right]\leq\frac{\big{(}2+K_{\text{conf}}\big{)}C_{h}}{\eta_{c}\sqrt{N}}+\frac{M_{c}}{\eta_{c}N},$
@@
with $K_{\text{mono}}\triangleq\Big{(}\big{(}2+K_{\text{conf}}\big{)}C_{h}+M_{c}\Big{)}/\eta_{c}$. ∎
F.3 Proof of Lemma 10 (Sufficient coverage)